Question 49 8 pts The isomerization of methylisocyanide, CH3NC - CH3CN, follows first-order kinetics. The half-lives...

The isomerization of methyl isocyanide, CH3NC - CH3CN, follows first-order kinetics. The half-lives were found to be 161 min at 199°C and 12.5 min at 230°C. Calculate the activation energy for this reaction. (R-8.314 J/mol K) 

  • 124 kJ/mol 

  • 6.17 x 10-3kJ/mol 

  • 78.2 kJ/mol 

  • 163 kJ/mol 

  • 31.4 kJ/mo
Answer #2


on Pation id K= A eki is rate constant K₂ is rate constant at temperature to at temperature T2 Ea is activation energy A is frequency factor kiKL= 0.693 0-693 161 min 0.004304 min fya K2 = 0.693 0-693 12.5 min O. 05544 min 0.004302 gmin 0.05544 min 412.15 Fairt-0.000
Answer #1

* using Arrhenius Equation in e in) = -fan het Taz 199+ 273 = 472 K T2 = = 230 +273 = 503K - Ea in 1925) 14 442) I 472 8.314